OpenStudy (anonymous):

Divide. Write a mixed numeral for the answer whenever possible. -42÷ 5 1/4

OpenStudy (anonymous):

I got -8. Is this right

OpenStudy (shamil98):

-42 x 4/21 = -168/21 = -8 yep you are right

OpenStudy (anonymous):

ye shams right so you're right

hartnn (hartnn):

yeah batman's right that sham's right that you're right :)

OpenStudy (yttrium):

yEAH @hartnn is right that @iambatman is right that @shamil98 is right that you are right. SUPER SUWAAAG!!

hartnn (hartnn):

lol, "right-ception" :P